### v1

@@ III-B Explicit World Model Construction
The surface geometries of the real-world object are usually complex, thus the extracted meshes can be non-convex. We further perform convex decomposition using VHACD [52] before importing the meshes to the physics simulator, which is essential for realistic physics simulation of robot interaction.
@@ III-C Sampling-based Model Predictive Control
Having an explicit physics model and a target joint state $s_{target}$ of the articulated object, the agent needs to search for a trajectory that can change the current joint state $s_{initial}=s_{1}$ to $s_{target}$. The expected relative joint movement is $\Delta s_{target}=s_{target}-s_{initial}$. Because of the complex contact between the robot end-effector and the constructed object model, the informative gradient of the objective function can hardly be acquired. Therefore, we employ sampling-based model predictive control, which is a zeroth-order method, to search for an optimal trajectory. There are various kinds of sampling-based model predictive control algorithms according to the zeroth-order optimization method used, such as Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [53], Cross-Entropy Method (CEM) [54], and Model Predictive Path Integral Control (MPPI) [55]. Among these methods, we select the iCEM method [56] to search for a feasible long-horizon trajectory to complete the task due to its simplicity and effectiveness. We briefly describe how we apply the iCEM method in the following paragraph.
@@
Trajectory length $T\in\mathbb{N}^{+}$ denotes the maximum time steps in a trajectory. At each time step $t(t<T)$, the action of the robot $\bm{a}_{t}\in\mathbb{R}^{d}$ is the incremental value of the joint position, where $d$ is the number of degrees of freedom (DOF) of the robot. The population $N$ denotes the number of samples sampled in each CEM iteration. Planning horizon $h$ determines the number of time steps the robot plans in the future at each time step. The top $K$ samples according to rewards compose an elite set, which is used to fit means and variances of a new Gaussian distribution. Please refer to [56] for details of the algorithm.
@@
At each time step $t$, the agent generates an action for the robot $\bm{a}_{t}\in\mathbb{R}^{d}$, where $d$ is the dimension of the action space. For 2-finger gripper tasks, $d=8$, which consists of the 7 DOF of the robot arm and the 1 DOF of the gripper. However, for dexterous hands, $d=23$, which includes the 7 DOF of the robot arm and the 16 DOF for the hand. The computational cost of iCEM is multiplied due to the high dimensionality of the action space. Consequently, directly searching in the original joint space of the multi-finger dexterous hand is not feasible. Moreover, the high-dimensional space of the dexterous hand may lead to unnatural postures. Therefore, it becomes essential to reduce the action space within the iCEM algorithm when using the dexterous hands.
@@
In our work, we propose to use eigengrasp [9] to reduce the high-dimensional action space associated with dexterous hands. This approach involves clustering a substantial number of grasping object actions to extract low-dimensional principal components. These components are then linearly combined to approximate the hand’s grasping posture. We first generate a dataset containing a diverse set of distinct grasping postures of the dexterous multi-finger hand in simulation using DexGraspNet [57]. We perform Principal Components Analysis(PCA) method on the dataset to get the largest $m$ eigenvectors $\boldsymbol{e}_{1}$, …, $\boldsymbol{e}_{m}$. The iCEM algorithm is then performed on the action space $\bm{a}_{t}\in\mathbb{R}^{7+m}$. The joint angles of the hand $\bm{q}_{h}$ are computed as a linear combination of the $m$ eigenvectors:
@@
III-C1 Two-finger gripper
For the two-finger gripper, the reward function consists of the following terms:
@@ III-C1 Two-finger gripper
$r_{target}=-\omega_{t}*(s_{target}-s_{t})/(s_{target}-s_{initial})$
This reward encourages $s_{t}$ to converge to $s_{target}$.
@@
This reward encourages the robot to have first contact with the object in the correct direction and to keep in contact with the object when moving the part. Also, this reward tries to prevent parts other than the fingertip or the target part of the object from colliding.
@@
$r_{dist}=\omega_{d}*||\bm{p}_{part}-\bm{p}_{grasp}||^{2}$
@@
$r_{reg}=-\sum_{i=0}^{d}{(\omega_{a}*a_{i}+\omega_{v}*v_{i})}$
@@
III-C2 Dexterous hand
For the dexterous hand, apart from the success reward $r_{success}$ and approaching reward $r_{target}$, which remain consistent in the 2-finger gripper’s reward function, the other three terms are as follows:
@@ III-C2 Dexterous hand
$r_{dist}=\omega_{d}*||\bm{p}_{part}-\bm{p}_{grasp}||^{2}$
This reward encourages the dexterous hand to get closer to the target part of the object, where $\bm{p}_{part},\bm{p}_{grasp}\in\mathbb{R}^{3}$ denotes the position of the geometry center of the target part and the grasp center of dexterous hand in Cartesian space, respectively.
@@
$r_{reg}=-\sum_{i=0}^{d}{(\omega_{v}*v_{i}+\omega_{p}*e_{p})}$
@@ IV EXPERIMENTS
In this section, we evaluate the precision and effectiveness of the proposed method for manipulating articulated objects for both two-finger grippers and dexterous hands. We first conduct a large number of real-world articulated object manipulation experiments and quantitatively compare the performance. Then we design 4 ablation studies to verify the effectiveness of different modules of our method. Finally, we validate the operational advantage of the dexterous hand against the two-finger gripper by comparing the task execution efficiency in simulation.
@@ IV-A Experimental Setup
Fig. 4 shows the real-world experimental setup. For the robot, a 7-DOF robot arm (ROKAE xMate3Pro) is used and an RGBD camera (Intel RealSense D415) is set to capture the visual input. The robot arm base is fixed at the table. Two kinds of end effectors are used: a 1-DoF 2-finger gripper (Robotiq 2F-140) and a 16-DoF 4-finger dexterous hand (Allegro Hand).
We choose 3 categories of common articulated objects for experiments, which are drawers, faucets and laptops as shown in Fig. 4. For the drawer, we assume that only one part of the drawer requires to be operated if there is more than one movable part. Besides, we only consider the case that the handle of the faucet rotates in horizontal direction. The articulated object is randomly located on the table with its base link fixed, and $s_{0}$ is randomly set. We randomly select $\Delta s_{target}$ which does not exceed the joint limit and covers both directions of possible movement.
To remove the influence of the background, we crop the object point cloud out of the scene using a bounding box. It is worth noting that we locate the camera on the right side of the robot rather than the front. This setting is better aligned with real application scenarios while increasing point cloud occlusion and manipulation difficulty. We further build the robot in simulation using the CAD models. We use SAPIEN [51] as the physics simulator to collect training data for the Explicit Physics Model Construction module and create simulation environments for the Sampling-based Model Predictive Control module.
Figure 4: (a) (b) Real-world experimental setup of two-finger gripper and dexterous hand. (c) Articulated object for manipulation. There are 3 categories of articulated objects in our real-world experiment.
IV-B Data Collection and Training
Explicit Physics Model Construction. For drawers and faucets, we choose 14 and 8 objects from Shape2Motion datasets [58]. For laptops, we choose 5 objects from PartNet-Mobility dataset [59] because the joint limits of laptops in PartNet-Mobility dataset are more reasonable than in Shape2Motion datasets. The original meshes are not water-tight in the PartNet-Mobility dataset, which cannot be used to compute point occupancy, so we use ManifoldPlus [60] to fix the meshes. We reimplement the data collection code using SAPIEN simulator to keep consistency. When collecting data, the object is fixed on the origin and then randomly rotated around the z-axis by $[-60^{\circ},+60^{\circ})$. The camera is on a sphere centered on the object’s center. Azimuth and Altitude of the camera are randomly sampled from $[-60^{\circ},+60^{\circ})$ and $[15^{\circ},45^{\circ})$. 10000 samples are collected for each category. We downsample the object point clouds to 8192 points. The 3 categories are trained jointly.
@@ IV-B Data Collection and Training
Eigengrasp Dataset Construction. To build the dataset for eigengrasp computation, we utilize DexGraspNet [57] to generate a collection of random grasping postures for the Allegro Hand. The dataset includes 60800 grasp postures across 474 objects. We then compute the eigengrasp based on this data. Fig. 5 shows the accumulated ratios of different eigengrasp dimensions. Unless otherwise specified, we use eigengrasp dimension $m=2$ for dexterous manipulation experiments.
@@
Figure 5: Results of accumulated ratios of different eigengrasp dimensions.
IV-C Experiments on 2-Finger Gripper
IV-C1 Real World Articulated Object Manipulation
@@
For parameters of the Interactive Perception module, we choose $n_{p}=10$ and $n_{a}=10$. For parameters of the Sampling-based Model Predictive Control module, we find that $T=50$, $N=300$, $h=10$ and $K=20$ are able to complete all the tasks. The range of incremental value of joint position is set to $[-0.05,0.05]$. The parameters in the reward function are determined manually according to experience in the simulation environment. We set $\omega_{s}=20$, $\epsilon=0.005$(m or rad), $\omega_{t}=50$, $\omega_{contact}=10$, $\omega_{collision}=60$, $\omega_{d}=10$, $\omega_{a}=0.01$ and $\omega_{v}=0.03$. We use 20 processes for sampling in simulation on a computer that has an Intel Core i7-12700 CPU and an NVIDIA 3080Ti GPU. It takes 4 minutes to find a feasible trajectory.
We conduct about 30 experiments for each category. After the trajectory is executed in the real world, we measure the real joint movement $\Delta s_{real}=s_{real}-s_{initial}$ and compare it with the target joint movement $\Delta s_{target}=s_{target}-s_{initial}$. We compute the error $\delta=\Delta s_{real}-\Delta s_{target}$ and the relative error $\delta_{r}={\delta}/{\Delta s_{target}}\times 100\%$, results of all the experiments can be found in Fig. 6, and statistical results can be found in Table I. Trajectories of both opening and closing the laptop are shown in Fig. 7.
Among all 3 categories, the drawer has the lowest $\left|\delta_{r}\right|$ and the faucet has the highest $\left|\delta_{r}\right|$ according to Table I. It is reasonable because the size of the faucet is relatively small, a minor inaccuracy in model construction or trajectory execution will result in a big error in the joint state. About $70\%$ of manipulations achieve a $\left|\delta_{r}\right|<30\%$ for drawers and laptops, which shows the accuracy of our method.
@@
TABLE I: Accuracy of real articulated objects manipulation with 2-finger gripper
Category Drawer Laptop Faucet
Number of manipulations 31 32 30
<10% 12 7 0
<30% 22 20 9
Number of manipulations s.t. $\left|\delta_{r}\right|$ <50% 28 26 19
Avg $\left|\delta\right|$ 1.15cm 5.69 10.37
Avg $\left|\delta_{r}\right|$ 21.81% 27.26% 56.21%
Errors may be caused by the following factors:
  1. (1)
    The constructed mesh is not accurate enough, especially for the parts that are occluded. For example, the inside face of the drawer front cannot be observed by the RGBD camera, so when the digital twin is constructed, the drawer front is thicker than the real one. It causes the results of opening tasks of drawers (which has average $\left|\delta\right|$ of $2\mathrm{cm}$) to be worse than closing tasks (which have average $\left|\delta\right|$ of $0.5\mathrm{cm}$). It is worth noting that there is a relative error of over $400\%$ in turning faucet tasks. This happens because the robot touches the part close to the joint axis first (which does not occur in the simulation), causing a huge rotation of the handle.
  2. (2)
    The dynamic properties of the real articulated objects are complicated. For example, the elastic deformation of laptops is not modeled in the simulation.
  3. (3)
    The kinematic structure of a real articulated object is not ideal. For example, there might be gaps in the drawer rails, which turns the original prismatic joint into a joint with several DOFs.
@@
IV-C2 Ablation Study on Reward Function
The reward function in the sampling-based model predictive control module is designed to guide the robot to complete the task. To examine the impact of each term of the reward function, we conduct the ablation study. There are 5 terms in the reward function, so 6 groups of experiments are conducted to reveal each term’s influence against the full reward function. The first group runs iCEM with the full reward function as in Section III-C. Each of the other 5 groups drops one term of the full reward function. In each group, 5 tasks are conducted to make the results more general. The task that is considered to be failed if not completed within 50 time steps. Fig. 8 summarizes the experimental results.

### v2

@@ III-B4 Kinematic Structure Construction
Finally, we transform the constructed model into the Unified Robot Description Format (URDF), which can be easily loaded into widely used multi-body physics simulators, such as SAPIEN [68]. Since the geometries of the real-world object are usually complex, the reconstructed meshes can be non-convex. We further perform convex decomposition using VHACD [69] before importing the meshes to the physics simulator, which is essential for realistic physics simulation of robot interaction.
@@ III-C Sampling-based Model Predictive Control
Having an explicit physics model and a target joint state $s_{target}$ of the articulated object, the agent needs to search for a trajectory that can change the current joint state $s_{initial}=s_{1}$ to $s_{target}$. The expected relative joint movement is $\Delta s_{target}=s_{target}-s_{initial}$. Because of the complex contact between the robot end effector and the constructed object model, the informative gradient of the objective function can hardly be acquired. Therefore, we employ sampling-based model predictive control, which is a zeroth-order method, to search for an optimal trajectory. There are various kinds of sampling-based model predictive control algorithms according to the zeroth-order optimization method used, such as Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [70], Cross-Entropy Method (CEM) [71], and Model Predictive Path Integral Control (MPPI) [72]. Among these methods, we select the iCEM method [73] to search for a feasible long-horizon trajectory to complete the task due to its simplicity and effectiveness. We briefly describe how we apply the iCEM method in the following paragraph.
@@
Trajectory length $T\in\mathbb{N}^{+}$ denotes the maximum time steps in a trajectory. At each time step $t(t<T)$, the action of the robot $\bm{a}_{t}\in\mathbb{R}^{d}$ is the incremental value of the joint position, where $d$ is the number of degrees of freedom (DOF) of the robot. The population denotes the number of samples sampled in each CEM iteration. Planning horizon $h$ determines the number of time steps the robot plans in the future at each time step. The top $E$ samples according to rewards compose an elite set, which is used to fit means and variances of a new Gaussian distribution. Please refer to [73] for details of the algorithm.
@@
At each time step $t$, the agent generates an action for the robot $\bm{a}_{t}\in\mathbb{R}^{d}$, where $d$ is the dimension of the action space. For suction grippers, $d=7$, which is the DOF of the robot arm. For 2-finger gripper tasks, $d=8$, which consists of the DOF of the robot arm and the 1 DOF of the gripper. However, for dexterous hands, $d=7+d_{\text{hand}}$, which includes the 7 DOF of the robot arm and the $d_{\text{hand}}$ DOF for the hand. The computational cost of iCEM is multiplied due to the high dimensionality of the action space. Consequently, directly searching in the original joint space of the multi-finger dexterous hand is not feasible. Moreover, the high-dimensional space of the dexterous hand may lead to unnatural postures. Therefore, it becomes essential to reduce the action space within the iCEM algorithm when using the dexterous hands.
@@
In our work, we propose to use eigengrasp [10] to reduce the high-dimensional action space associated with dexterous hands. This approach involves clustering a substantial number of grasping object actions to extract low-dimensional principal components. These components are then linearly combined to approximate the hand’s grasping posture. We first generate a dataset containing a diverse set of distinct grasping postures of the dexterous multi-finger hand in simulation using DexGraspNet [74]. We perform Principal Components Analysis (PCA) on the dataset to obtain the first $m$ eigenvectors $\boldsymbol{e}_{1}$, …, $\boldsymbol{e}_{m}$ with largest eigenvalue. The iCEM algorithm is then performed on the action space $\bm{a}_{t}\in\mathbb{R}^{7+m}$. The joint angles of the hand $\bm{q}_{h}$ are computed as a linear combination of the $m$ eigenvectors:
@@
III-C1 Suction gripper
For the suction gripper, the reward function consists of the following terms:
@@ III-C1 Suction gripper
$r_{target}=-\omega_{t}\cdot\frac{s_{target}-s_{t}}{s_{target}-s_{initial}}$
This reward encourages $s_{t}$ to converge toward $s_{target}$.
@@
This reward encourages the robot to establish initial contact with the object in the correct direction and maintain contact while moving the part. Additionally, it prevents collisions of parts other than the fingertip or the target part of the object.
@@
$r_{dist}=\omega_{d}\cdot||\bm{p}_{virtual}-\bm{p}_{grasp}||_{2}^{2}$
It is important to note that, unlike the distance reward for two-finger grippers or dexterous hands, which encourage the entire end effector to approach the target link of the object, the suction gripper focuses on bringing the tip of the suction closer to the target link. To achieve this, we mount a virtual link, $\bm{l}_{virtual}$, at the tip of the suction gripper for reward computation. This reward encourages the suction gripper to move closer to the target part of the object, where $\bm{p}_{grasp},\bm{p}_{virtual}\in\mathbb{R}^{3}$ represent the position of the geometry center of the target part and the position of the suction tip, respectively.
(5) access direction reward
$r_{dir}=\begin{cases}\omega_{dir},\quad\text{if}\quad\dfrac{\mathbf{v}_{suction}\cdot\mathbf{n}_{target}}{||\mathbf{v}_{suction}||_{2}\cdot||\mathbf{n}_{target}||_{2}}\geq\cos\psi\\ 0,\qquad\qquad\text{else}\end{cases}$
This reward function encourages the suction to make contact with the target link’s surface within a certain range of directions. $\mathbf{v}_{suction}$ and $\mathbf{n}_{target}$ represent the orientation of the suction gripper and the normal of the target link’s surface, respectively. $\psi$ is a predefined angle threshold.
III-C2 Two-finger gripper
For the two-finger gripper, apart from success reward $r_{success}$, approaching reward $r_{target}$ and contact reward $r_{contact}$ which remain consistent in the reward function of suction gripper, the other two terms are as follows:
(1) distance reward
$r_{dist}=\omega_{d}\cdot||\bm{p}_{part}-\bm{p}_{grasp}||_{2}^{2}$
@@ III-C2 Two-finger gripper
$r_{reg}=-\sum_{i=1}^{d}{(\omega_{a}\cdot a_{i}+\omega_{v}\cdot v_{i})}$
@@
III-C3 Dexterous hand
For the dexterous hand, apart from the success reward $r_{success}$ and approaching reward $r_{target}$, which remain consistent in the suction gripper’s reward function, the other three terms are as follows:
@@ III-C3 Dexterous hand
$r_{dist}=\omega_{d}\cdot||\bm{p}_{part}-\bm{p}_{grasp}||_{2}^{2}$
This reward encourages the dexterous hand to get closer to the target part of the object, where $\bm{p}_{part},\bm{p}_{grasp}\in\mathbb{R}^{3}$ denotes the position of the geometry center of the target part and the grasp center of dexterous hand, respectively.
@@
$r_{reg}=-\omega_{p}\cdot e_{p}-\sum_{i=1}^{d}{(\omega_{v}\cdot v_{i})}$
@@ IV EXPERIMENTS
In this section, we evaluate the precision and effectiveness of the proposed method for manipulating articulated objects using suction grippers, two-finger grippers and dexterous hands. We first conduct a large number of real-world articulated object manipulation experiments and quantitatively compare the performance. Then we design ablation studies to verify the effectiveness of different modules of our method, and validate the operational advantage of the dexterous hand against the two-finger gripper by comparing the task execution efficiency in simulation. Finally, we analyze the explicit world model construction accuracy and its impact on manipulation performance.
@@ IV-A Experimental Setup
Fig. 5 shows the real-world experimental setup. For the robot, a 7-DOF robot arm (ROKAE xMate3Pro) is used, which is fixed at the table. An RGBD camera (Intel RealSense D415) is set to capture the visual input and an additional RGB camera is positioned at a side viewpoint to capture side images that assist explicit model construction. Note that only the RGBD camera’s intrinsic parameters and transformation relative to the robot base need to be calibrated, whereas no calibration is required for the RGB camera. 4 kinds of end effectors are used: a suction gripper, a 1-DoF 2-finger gripper (Robotiq 2F-140), a 16-DoF 4-finger dexterous hand (Allegro Hand), a 12-DoF 5-finger dexterous hand (XHand).
We choose 4 categories of common articulated objects for experiments, which are drawers, cabinets, lamps, and laptops as shown in Fig. 6.
The articulated object is randomly located on the table with its base link fixed, and $s_{0}$ is randomly set. We randomly select the goal state $\Delta s_{target}$ which does not exceed the joint limit and covers both directions of possible movement. We use a double-arm protractor ruler to measure the object’s state after manipulation and compute the errors.
We use SAPIEN [68] as the physics simulator to collect training data for the Explicit Physics Model Construction module and create simulation environments for the Sampling-based Model Predictive Control module.
Figure 5: Real-world experimental setup of (a) suction gripper; (b) two-finger gripper and (c)(d)dexterous hands. Here we use another RealSense D415 as the Side camera.
Figure 6: Articulated objects for manipulation.
IV-B Implementation Details
For affordance learning from self-supervised interaction, we introduce modifications to Where2Act [16] to accommodate our experimental setting: (1) we randomly set the target joint state; (2) the azimuth of the camera is randomly sampled from $[-60^{\circ},+60^{\circ})$ and the altitude is randomly sampled from $[15^{\circ},45^{\circ})$; (3) we move the center of the point cloud to the origin. We downsample the object point cloud to 2000 points before feeding it into the network. For real-world deployment, we centralize the captured point cloud using an estimated bounding box.
For the 3D post-contact vector generation of VRB, we position the virtual camera at $(0.2,0,0.7)$ in the robot base frame, with a pitch angle of $\theta=52^{\circ}$. This configuration ensures a balanced between preserving visual features for the affordance network and ensuring sufficient offset for accurate computation of the 3D post-contact vector.
@@ IV-B Implementation Details
To build the dataset for eigengrasp computation, we utilize DexGraspNet [74] to generate a collection of random grasping postures for the Allegro Hand and the XHand. The dataset includes 60800 grasp postures across 474 objects. We then compute the eigengrasp based on this data. Fig. 7 shows the accumulated ratios of different eigengrasp dimensions. Interestingly, we find that the trends of the two hands, despite having different numbers of DoFs, are well aligned, suggesting that the eigengrasp effectively captures the fundamental features of the hand posture distribution. Unless otherwise specified, we use eigengrasp dimension $m=2$ for dexterous manipulation experiments.
@@
Figure 7: Results of accumulated ratios of different eigengrasp dimensions for Allegro Hand and XHand.
IV-C Experiments on Suction Gripper
IV-C1 Implementation Details
@@
For parameters of the Sampling-based Model Predictive Control module, $T=50$, $E=20$, and $h=10$ are able to complete all the tasks. The range of incremental value of joint position is set to $[-0.05,0.05]$. The parameters in the reward function are determined manually according to experience in the simulation environment. We set $\epsilon=0.005$ for prismatic joints, $\epsilon=0.02$ for revolute joints, $\omega_{s}=20$, $\omega_{t}=50$, $\omega_{contact}=10$, $\omega_{collision}=60$, $\omega_{d}=10$, $\omega_{a}=0.01$, $\omega_{v}=0.03$ and $\psi=15^{\circ}$. We use 30 processes for sampling in simulation on a computer that has an AMD Ryzen 9 9950X 16-Core CPU and an NVIDIA 3080Ti GPU. It takes about 1.5 minutes to find a feasible trajectory.
IV-C2 Comparison With Baseline
We choose UMPNet [22] as the baseline method, as it can also achieve goal-conditioned manipulation. We choose two categories: Laptop and Cabinet and conduct 10 experiments for each category. After the trajectory is executed in the real world, we measure the real joint movement $\Delta s_{real}=s_{real}-s_{initial}$ and compare it with the target joint movement $\Delta s_{target}=s_{target}-s_{initial}$. We compute the error $\delta=\Delta s_{real}-\Delta s_{target}$ and the relative error $\delta_{r}={\delta}/{\Delta s_{target}}\times 100\%$. Results of all the experiments can be found in Fig. 8, and statistical results can be found in Table I. As shown, our method achieves significantly lower error compared to UMPNet. This improvement may be attributed to the fact that UMPNet uses a fixed step size, whereas our method employs continuous values in MPC. Additionally, the two images representing the current state and the goal state may not provide sufficient information for precise manipulation in UMPNet.
Figure 8: Comparison of our method and UMPNet on real-world articulated object manipulation with suction gripper.
TABLE I: Accuracy comparison of real articulated objects manipulation with suction gripper
Category Laptop Cabinet
Method UMPNet [22] Ours UMPNet [22] Ours
[…]
Avg $\left|\delta\right|$downarrow 16.0 7.2 15.3 6.8
Avg $\left|\delta_{r}\right|$$\downarrow$ 42.9% 23.3% 41.5% 21.7%
IV-D Experiments on 2-Finger Gripper
IV-D1 Real World Articulated Object Manipulation
For parameters of the Interactive Perception module, we choose $n_{p}=10$ and $n_{a}=10$. For parameters of the Sampling-based Model Predictive Control module, we find that $T=50$, $E=300$, and $h=10$ are able to complete all the tasks. The range of incremental value of joint position is set to $[-0.05,0.05]$. The parameters in the reward function are determined manually according to experience in the simulation environment. We set $\omega_{s}=20$, $\epsilon=0.005$ (m or rad), $\omega_{t}=50$, $\omega_{contact}=10$, $\omega_{collision}=60$, $\omega_{d}=10$, $\omega_{a}=0.01$ and $\omega_{v}=0.03$. We use 30 processes for sampling in simulation on a computer that has an AMD Ryzen 9 9950X 16-Core CPU and an NVIDIA 3080Ti GPU. It takes about 2.5 minutes to find a feasible trajectory.
We choose two categories: Laptop and Cabinet, and conduct 20 experiments for each category. Results of all the experiments can be found in Fig. 9, and statistical results can be found in Table II. Trajectories of both opening and closing the laptop are shown in Fig. 10. About $75\%$ of manipulations achieve a $\left|\delta_{r}\right|<30\%$, which shows the accuracy of our method.
@@ IV-D1 Real World Articulated Object Manipulation
TABLE II: Accuracy of real articulated objects manipulation with 2-finger gripper
Category Laptop Cabinet
Number of manipulations 20 20
<10% 6 4
<30% 16 15
Number of manipulations s.t. $\left|\delta_{r}\right|$ <50% 20 18
Avg $\left|\delta\right|$ 7.2 7.4
Avg $\left|\delta_{r}\right|$ 21.5% 22.5%
IV-D2 Ablation Study on Reward Function
The reward function in the sampling-based model predictive control module is designed to guide the robot to complete the task. To examine the impact of each term of the reward function, we conduct the ablation study. There are 5 terms in the reward function, so 6 groups of experiments are conducted to reveal each term’s influence against the full reward function. The first group runs iCEM with the full reward function as in Section III-C. Each of the other 5 groups drops one term of the full reward function. In each group, 4 tasks are performed to make the results more general. The task that is considered to be failed if not completed within 50 time steps. Fig. 11 summarizes the experimental results.